# Towards Exploring Continual Fine-Tuning for Enhancing Language Ability in Large Language Models

### v1

@@ 4 Mitigating Strategies for Multilingual CFT
To mitigate the decline in task ability, we study two CFT techniques, Generative Replay (GR) and heuristic-based Layer Freezing (LF). In Generative Replay, we consider a new English data generation method motivated by the correlation between dataset similarity and task ability (§3.2). With heuristic-based Layer Freezing, we employ specific heuristics to find out the subset of layers to freeze in the model during Phase 2 fine-tuning.
4.1 Generative Replay
@@
Typically, Generative Replay (GR) is a technique that generates data from past distributions to be used alongside new task data for the continual fine-tuning of a model on a new task (shin2017continuallearningdeepgenerative). However, from §3.2, we see that if the phase-wise datasets encode similar tasks, the decline in task ability is mitigated. Based on this observation, we use the Phase 1 model to generate responses, in English, from the English counterpart of the multilingual dataset used for training in Phase 2. This generated replay dataset acts as a bridge between the distributions of Phase 1 and Phase 2.
@@
During Phase 2 fine-tuning, we include varying quantities of this generated data: specifically, 5% (GR_5) and 10% (GR_10), of the Phase 2 dataset. As a baseline, we also fine-tune the models with a similar sized subset of the English counterpart with original responses [^6] This dataset may not be available for all multilingual datasets eg. Aya (singh2024ayadatasetopenaccesscollection). In that case, instructions can always be translated to English but it is not always practical to translate responses. Hence, this baseline is the best-case scenario for our GR strategy.. We refer to this baseline as English Replay (ER_10).
4.2 Heuristic-based Layer Freezing
@@
Model regularization is an effective technique to mitigate the drop in the previous task’s performance in continual learning (e.g., EWC (kirkpatrick2017overcoming)). However, this is computationally inefficient as it requires using both the old and new sets of parameters. Instead, we use Layer Freezing (LF), a relatively efficient technique for use as a ‘regularizer’ to preserve task ability during Phase 2. We consider the following two variations to select the set of layers to freeze:
@@
    LF_H2: freezing the top-10 layers that have changed the most during Phase 1 fine-tuning (e.g., Mistral-7B Base to Mistral-7B-Instruct). We select these layers separately for Key, Query, and Value, for each attention head.
We present our results in Table 3 for both GR and LF. Along with English Replay (ER), we define another baseline in which we use LoRA (hu2022lora) for continually fine-tuning in Phase 2.
@@ 4.3 Results Discussion
From Table 3, we see that GR and LF successfully mitigate the decline in task ability and also show gains in language ability. For instance, Mistral-7B with GR_5 achieves better performance in MLQA and XLSUM when fine-tuned with MultiAlpaca. We also close the gap with Mistral-7B-Instruct  on IFEval, Alpaca Eval, MMLU, and HellaSwag with our mitigation strategies.
(a) Mistral-7B
(b) LLaMA-3-8B
Figure 3: Demonstrating extent of cross-lingual transfer in Mistral-7B and LLaMA-3-8B on a parallel dataset prepared by subsampling Flores (costa2022no). We find that the English activation cluster for LLaMA-3-8B is separated from the multilingual cluster, compared to Mistral-7B.
LLaMA-3-8B Doesn’t Show Consistent Improvement with our Mitigations.
From Table 3, while both GR and LF improve on the baseline LLaMA-3-8B-Instruct MultiAlpaca, the gains in task and language ability are not comparable to LLaMA-3-8B-Instruct.
To understand this further, for GR, we investigate the cross-linguality difference between LLaMA-3-8B and Mistral-7B. Like Figure 1, we plot t-SNEs of the mean model activations for the Mistral-7B and LLaMA-3-8B base models on two parallel datasets, English and Multilingual. We create the parallel datasets by subsampling data from Flores (costa2022no). In Figure 3, we see that the English activation cluster for LLaMA-3-8B is separated out from multilingual cluster, compared to Mistral-7B. This suggests that GR may not be as effective when the model has less cross lingual ability. While for LF, we acknowledge that our method to identify the layers to freeze may not be the best and better methods to identify which layers to freeze can be a direction for future work.
With regards to LLaMA-3-8B-Instruct MultiAlpaca LA results in Table 2, we believe that this is due to lack of cross-linguality in LLaMA-3-8B-Instruct and less data in MultiAlpaca which fails to cause sufficient representation drift to improve the model’s performance.
Forgetting with LoRA. For Mistral-7B-Instruct and LoRA fine-tuning, we see an increase in language ability but a decline in task ability. But the decline is not as much as full fine-tuning. For LLaMA-3-8B-Instruct and LoRA, there is a greater decline in both task and language ability. The decline is similar (or slightly lower) than the full fine-tuning scenario. These results show that LoRA also suffers from forgetting when used for continual fine-tuning.
Additional Results. In the Appendix (§C), we repeat the same experiment from §3.4 to quantify the representation change in the fine-tuned models using our mitigating strategies. We see a trend similar to Figure 2. That is, a decrease in the variation in the model activations, compared to the baseline model trained on Instruct and MultiAlpaca. The decrease is more pronounced for Mistral-7B compared to LLaMA-3-8B. In Appendix §C, we also present TA and LA results for the Instruct-mOpenOrca phase-wise datasets.
@@ 5 Conclusion & Future Work
In this paper, to the best of our knowledge, we present a first study on the influence of the similarity of phase-wise datasets on the task and language adaptability of LLMs through CFT. Through extensive experiments on the Mistral-7B and LLaMA-3-8B models, we show that when datasets are similar, task ability is preserved; otherwise, it declines. To mitigate this, we propose layer freezing and generative replay as mitigating strategies based on specific heuristics. Our results indicate that these strategies help improve task performance while not compromising on the LLM’s language adaptability.
@@
Our results show that there is no one-size-fits-all strategy to mitigate decline in task ability, among the strategies discussed. Future work can explore developing other parameter-efficient regularization methods that address the current computational challenges with methods like EWC or forgetting due to LoRA. One can also explore analytical notions for task similarity in datasets.
@@ A.1 Hyperparameters for Fine-tuning and Training Setup
Hyperparameter Value Learning Rate $1\times 10^{-6}$ Epochs 4 Global Batch size 16 Scheduler Cosine Warmup Linear Warmup Steps 10 Optimizer AdamW (loshchilov2019decoupled) Weight Decay 0
Table A1: Hyperparameters for continual fine-tuning
@@ A.2 Fine-tuning Datasets
OpenOrca is an English-only self instruct dataset, created to best mimic the Orca dataset (mukherjee2023orca), which is not publicly available. To create the multilingual version of OpenOrca, namely mOpenOrca, we follow sphinx to generate selective translations for a subset of OpenOrca. The subset contains 50k samples from the OpenOrca dataset and we selectively translate them to 11 languages which are also in MultiAlpaca. In total, we generate 550k examples for all languages.
@@ A.3 Evaluation Tasks
    IFEval (zhou2023instruction): Instruction-Following Evaluation (IFEval) asses the ability of an LLM to follow natural language instructions. It comprises 500 verifiable instructions (e.g., “mention the keyword AI 3 times”). We choose IFEval as the instructions are verifiable and also test an LLM’s context understanding.
@@
    Alpaca Eval (li2023alpacaeval): This is an LLM-based automatic evaluator for instruction following models, to measure task ability. Like aggarwal2024maple, we evaluate our CFT models against text-davinci-003 responses on 800 instructions and use GPT4 (gpt-4-32k) as the evaluator.
@@
    MMLU (hendryckstest2021): Massive Multitask Language Understanding (MMLU) is a benchmark to assess an LLM’s knowledge and problem-solving abilities. It includes 57 subjects across domains like STEM, or law, with 16k MCQs in total.
@@
    HellaSwag (zellers2019hellaswag): This is a popular benchmark to evaluate the commonsense reasoning ability of an LLM. HellaSwag’s test split contains 10k samples in total.
Language Ability (LA).
@@
  • Question Answering: MLQA (lewis2019mlqa) contains 5k extractive question-answering instances in 7 languages. The XQuAD dataset (Artetxe:etal:2019) consists of a subset of 240 paragraphs and 1190 question-answer pairs across 11 languages.
@@
  • Summarisation: XLSUM (hasan-etal-2021-xl) spans 45 languages, and we evaluate our models in Arabic, Chinese-Simplified, English, French, Hindi, Japanese, and Spanish.
@@
Model Phase 1 (P1) Phase 2 (P2) IFEval ()($\uparrow$)( ↑ ) Alpaca Eval ()($\uparrow$)( ↑ ) MMLU ()($\uparrow$)( ↑ ) HellaSwag ()($\uparrow$)( ↑ ) Average Dataset Dataset P1 P2 P1 P2 P1 P2 P1 P2 P1 P2 Mistral-7B OpenOrca mOpenOrca 0.494 0.482 0.31 0.32 0.601 0.582 0.612 0.562 0.504 0.487 Instruct 0.550 0.426 0.35 0.06 0.575 0.507 0.641 0.509 0.529 0.376 LLaMA-3-8B OpenOrca 0.377 0.425 0.09 0.07 0.579 0.599 0.571 0.564 0.404 0.415 Instruct 0.735 0.205 0.14 0.0 0.340 0.236 0.533 0.250 0.437 0.173
@@
Table A1: Task Ability results for two-phase Continual Fine-tuning (CFT). With green, we highlight an increase in a model’s task ability post P2 fine-tuning. Likewise, red highlights a decline in a model’s task ability.
@@
Model Phase 1 Phase 2 MLQA ()($\uparrow$)( ↑ ) XLSUM ()($\uparrow$)( ↑ ) XQuAD ()($\uparrow$)( ↑ ) Average Dataset Dataset Phase 1 Phase 2 Phase 1 Phase 2 Phase 1 Phase 2 Phase 1 Phase 2 Mistral-7B OpenOrca mOpenOrca 0.435 0.36 0.007 0.008 0.556 0.643 0.332 0.337 Instruct 0.246 0.155 0.012 0.040 0.351 0.323 0.203 0.173 LLaMA-3-8B OpenOrca 0.401 0.453 0.017 0.006 0.499 0.531 0.306 0.330 Instruct 0.609 0.604 0.048 0.048 0.712 0.713 0.456 0.455
@@
Table A2: Language Ability results for two-phase Continual Fine-tuning (CFT). With green, we highlight an increase in a model’s language ability post Phase 2 fine-tuning. Likewise, red highlights a decline in a model’s language ability.
Appendix B Evaluating Language Ability for Multilingual Continual Fine-tuning
Task Ability.
@@ Appendix B Evaluating Language Ability for Multilingual Continual Fine-tuning
Table A1 present the task ability numbers of our ablations on the OpenOrca-mOpenOrcaand Instruct-mOpenOrcadatasets using Mistral-7B and LLaMA-3-8B models. When the datasets are pairwise not similar, i.e., Instruct-mOpenOrca, Mistral-7B shows a significant decline in the average task ability, from 0.529 in Phase 1 to 0.376 in Phase 2. Likewise, LLaMA-3-8B also experiences a decrease, dropping from 0.437 to 0.173 on average.
@@
In contrast, when the pairwise datasets are similar, i.e., OpenOrca and mOpenOrca, Mistral-7B sees a marginal drop between the phases $(0.504\rightarrow 0.487)$, on average. LLaMA-3-8B’s performance sees an improvement in the average task ability, from 0.404 to 0.415.
Language Ability.
@@
Table A2 tabulates the results for language ability. We see an improvement in the average language ability for the OpenOrca-mOpenOrca dataset pair, for both Mistral-7B and LLaMA-3-8B. For Instruct-mOpenOrca, with LLaMA-3-8B, the average language ability is virtually the same across tasks. However, for Mistral-7B, we see a slight drop in the average language ability, driven primarily due to a drop in performance for MLQA.
@@
Appendix C Mitigating Strategies

### v3

@@ 5 Mitigating Strategies for CFT
To mitigate EA decline, we explore two tailored CFT techniques [^7] In §F, we discuss how MAD-X Pfeiffer et al. (2020) and other PEFT-based methods Badola et al. (2023) are not suitable for our setting.: Distribution Replay and Layer Freezing. In Distribution Replay, we study Generative Replay (GR), a new English data generation method inspired by dataset similarity and English ability (§4.2), and English Replay (ER), which replays parallel English data of Phase 2’s distribution. In Layer Freezing (LF), we identify layers to freeze during Phase 2 fine-tuning using specific heuristics.
5.1 Distribution Replay
@@
Typically, Generative Replay (GR) is a technique that generates data from past distributions to be used alongside new task data for the continual fine-tuning of a model on a new task Shin et al. (2017). However, from §4.2, we do not see a decline in English ability if the phase-wise datasets encode similar instructions. Based on this, we use the Phase 1 model to generate responses, in English, from the English counterpart of the multilingual dataset used for fine-tuning in Phase 2. The intuition is that the generated dataset may bridge the distributions of Phase 1 and Phase 2.
@@
During Phase 2 fine-tuning, we include varying quantities of this generated data: specifically, 5% (GR_5) and 10% (GR_10), of the Phase 2 dataset. We also fine-tune the models with a similar sized subset of the English counterpart with original responses [^8] This dataset may not be available for all multilingual datasets, such as Aya Singh et al. (2024b). While instructions can be translated into English, translating responses is often impractical. Thus, ER is the best-case scenario for GR.. We refer to this mitigating strategy as English Replay (ER_10).
5.2 Layer Freezing
@@
Model regularization is an effective technique to mitigate the drop in the previous task’s performance in continual learning (e.g., EWC Kirkpatrick et al. (2017)). However, this is computationally inefficient as it requires using both the old and new sets of parameters. Instead, we use Layer Freezing (LF), a relatively efficient technique for use as a ‘regularizer’ to preserve English ability during Phase 2. We consider the following variations to select the set of layers to freeze. These variants allow us to study how different criteria for selecting frozen layers affect the trade-off between preserving English ability and enabling multilingual adaptation.
@@
    LF_H2: freezing the top-10 layers that have changed the most during Phase 1 fine-tuning (e.g., Mistral-7B Base to Mistral-7B-Instruct). We select layers separately for Key, Query, and Value, for each attention head.
  3. 3.
    Spectrum Hartford et al. (2024): freeze the "most informative" layers of the Phase 1 model based on their signal-to-noise ratio (refer to §D.1).
We present our results in Table 5 for both GR and LF. We define a baseline in which we use LoRA Hu et al. (2022) [^9] Parameter efficient techniques like LoRA Hu et al. (2022) are also widely used to efficiently fine-tune LLMs on multilingual data. However, such techniques also show forgetting on English Aggarwal et al. (2024) after Phase 2.for continually fine-tuning in Phase 2. We perform LoRA fine-tuning with rank 64 and quantisation bfloat16.
@@ 5.3 Results Discussion
From Table 5, we see that GR, ER, and LF mitigate the decline in EA and also show gains in MA.
Distribution Replay.
ER_10 demonstrates the best performance in both English and combined ability, with EA scores of 0.404 for Mistral-7B and 0.345 for LLaMA-3-8B, and the best combined average. GR_5 also excels in multilingual tasks, performing similar to ER_10: 0.298 vs. 0.276 for Mistral-7B and 0.348 vs. 0.359 for LLaMA-3-8B. GR_5 also performs reasonably well on English tasks, achieving scores of 0.394 and 0.236 for Mistral-7B and LLaMA-3-8B, respectively, making it a competitive strategy.
Layer Freezing.
Compared to ER and GR, LF_H1, LF_H2, and Spectrum show mixed results. LF_H2 performs better than LF_H1. Spectrum’s EA scores are better than LF_H1 and LF_H2, but suffers from lower multilingual numbers.
Additional Discussion & Results.
In §D.3, we also present EA and MA results for Mistral-7B Instruct-mOpenOrca for our mitigating strategies. Here, LF, particularly Spectrum, performs better than the other strategies. Furthermore, in §D.4, we analyze the computational cost of these strategies over the baseline CFT setup.
5.4 Controlling Representation Drift under Mitigation
Figure 3: Layer-wise covariance shift under mitigation strategies in Phase 2 CFT. We plot $\|\Sigma_{\text{Phase 2}}-\Sigma_{\text{Phase 1}}\|_{2}$ across layers. The vanilla Phase 2 model (Figure 2) exhibits a sharp increase in drift in higher layers. Replay (GR) substantially reduces the magnitude of this drift across all layers, while freezing (LF_H2) confines it to a smaller subset of layers.
Building on Figure 2, we examine how mitigation strategies affect representation drift relative to the vanilla Phase 2 baseline (Mistral-7B Instruct $\rightarrow$ MultiAlpaca). As in §4.4, we measure drift as $\|\Sigma_{\text{Phase 2}}-\Sigma_{\text{Phase 1}}\|_{2}$, where larger values indicate greater deviation from Phase 1 representations.
Consistent with earlier observations, the mitigation strategies explicitly designed to curb representational change (LF & GR) exhibit lower drift than the baseline model across layers. Replay-based methods (GR) reduce the overall magnitude of drift while preserving its layer-wise structure, whereas freezing-based methods (LF_H2) constrain the drift to a subset of layers.
Notably, generative replay remains closest to the Phase 1 representation geometry among all strategies. This closer alignment correlates with improved task and language performance compared to the vanilla Phase 2 model (Table 1 and Table 2). These results reinforce our central hypothesis: in two-phase CFT with dissimilar datasets, representation shift is the primary driver of English degradation, and explicitly controlling this shift preserves alignment with Phase 1 representations, leading to improved English retention without sacrificing multilingual gains.
@@ 6 Conclusion & Future Work
We study the role of dataset similarity in a novel two-phase continual fine-tuning (CFT) for multilingual adaptation. Across Mistral-7B and LLaMA-3-8B, we show that alignment between Phase 1 and Phase 2 datasets is critical: similar datasets preserve English ability while improving multilingual performance, whereas misaligned datasets lead to degradation. We identify representation shift as the underlying mechanism driving this behavior. In particular, misaligned fine-tuning induces large covariance shifts in hidden representations, concentrated in higher layers. We show that mitigation strategies such as generative replay and layer freezing improve English retention by controlling this drift, either by globally reducing its magnitude or by constraining it to specific layers, while preserving multilingual gains.
@@
Our results suggest several directions for future research. First, designing adaptive mitigation strategies that dynamically control representation drift during training could provide stronger and more efficient trade-offs than fixed heuristics Zhang et al. (2026). Second, developing principled measures of instruction-level dataset similarity may enable better prediction and prevention of degradation. Finally, exploring parameter-efficient approaches that explicitly regularize representation geometry could reduce the computational overhead of current methods while maintaining their effectiveness.
7 Limitations
Our study has several limitations. First, we rely on DES and MPD as proxies for dataset similarity; while effective in our setting, these metrics may not capture all nuances of instruction-level similarity. Second, our experiments are limited to Mistral-7B and LLaMA-3-8B, and the observed trends may not fully generalize to models with different architectures, scales, or training paradigms. Similarly, the effectiveness of our mitigation strategies (i.e., replay and layer freezing) may vary across datasets and model families. Third, our experimental setup is constrained by computational resources. We were unable to explore larger models, longer training schedules, or broader hyperparameter sweeps, which may further influence the observed trade-offs. Fourth, some mitigation strategies impose additional requirements. In particular, the best-performing method, ER_10, assumes access to parallel data, which may not always be available in practice. Finally, our evaluation focuses on specific benchmarks for task and language ability. While these provide controlled comparisons, they may not fully capture real-world performance across diverse applications.
@@ A.1 Hyperparameters for Fine-tuning and Training Setup
Hyperparameter Value Learning Rate $1\times 10^{-6}$ Epochs 4 Global Batch size 16 Scheduler Cosine Warmup Linear Warmup Steps 10 Optimizer AdamW Loshchilov and Hutter (2019) Weight Decay 0
Table A1: Hyperparameters for our Two-phase Continual Fine-tuning (CFT)
@@ A.2 Fine-tuning Datasets
OpenOrca is an English-only self instruct dataset, created to best mimic the Orca dataset Mukherjee et al. (2023), which is not publicly available. To create the multilingual version of OpenOrca, namely mOpenOrca, we follow Ahuja et al. (2024b) to generate selective translations for a subset of OpenOrca. The subset contains 50k samples from the OpenOrca dataset and we selectively translate them to 11 languages which are also in MultiAlpaca. In total, we generate 550k examples for all languages.
@@ A.3 Evaluation Tasks
    IFEval Zhou et al. (2023): Instruction-Following Evaluation (IFEval) asses the ability of an LLM to follow natural language instructions. It comprises 500 verifiable instructions (e.g., “mention the keyword AI 3 times”). We choose IFEval as the instructions are verifiable and also test an LLM’s context understanding.
@@
    Alpaca Eval Li et al. (2023): This is an LLM-based automatic evaluator for instruction following models, to measure task ability. Like Aggarwal et al. (2024), we evaluate our CFT models against text-davinci-003 responses on 800 instructions and use GPT4 (gpt-4-32k) as the evaluator.
@@
    MMLU Hendrycks et al. (2021): Massive Multitask Language Understanding (MMLU) is a benchmark to assess an LLM’s knowledge and problem-solving abilities. It includes 57 subjects across domains like STEM, or law, with 16k MCQs in total.
@@
    HellaSwag Zellers et al. (2019): This is a popular benchmark to evaluate the commonsense reasoning ability of an LLM. HellaSwag’s test split contains 10k samples in total.
Multilingual Ability (MA).
@@
  • Question Answering: MLQA Lewis et al. (2019) contains 5k extractive question-answering instances in 7 languages. The XQuAD dataset Artetxe et al. (2019) consists of a subset of 240 paragraphs and 1190 question-answer pairs across $1$ languages.
@@
  • Summarisation: XLSUM Hasan et al. (2021) spans 45 languages, and we evaluate our models in Arabic, Chinese-Simplified, English, French, Hindi, Japanese, and Spanish.
Appendix B Evaluating Multilingual Ability for Continual Fine-tuning
@@
Two-phase Continual Fine-tuning Model Phase 1 (P1) Phase 2 (P2) IFEval $(\uparrow)$ Alpaca Eval $(\uparrow)$ MMLU $(\uparrow)$ HellaSwag $(\uparrow)$ XLSUM_en $(\uparrow)$ Average Dataset Dataset P1 P2 P1 P2 P1 P2 P1 P2 P1 P2 P1 P2 Mistral-7B OpenOrca mOpenOrca 0.494 0.482 0.31 0.32 0.601 0.582 0.612 0.562 0.02 0.12 0.407 0.413 Instruct mOpenOrca 0.550 0.426 0.35 0.06 0.575 0.507 0.641 0.509 0.13 0.0 0.450 0.30 LLaMA-3-8B OpenOrca mOpenOrca 0.377 0.425 0.09 0.07 0.579 0.599 0.571 0.564 0.07 0.12 0.337 0.356 Instruct mOpenOrca 0.735 0.205 0.14 0.0 0.340 0.236 0.533 0.250 0.11 0.0 0.372 0.138 Dataset Mixture Model Dataset Mixture IFEval $(\uparrow)$ Alpaca Eval $(\uparrow)$ MMLU $(\uparrow)$ HellaSwag $(\uparrow)$ XLSUM_en $(\uparrow)$ Average Mistral-7B OpenOrca mOpenOrca 0.228 0.035 0.284 0.444 0.02 0.202 LLaMA-3-8B OpenOrca mOpenOrca 0.248 0.072 0.484 0.473 0.0 0.255
@@
Table B2: English Ability results for two-phase Continual Fine-tuning (CFT). With green, we highlight an increase in a model’s task ability post P2 fine-tuning. Likewise, red highlights a decline in a model’s task ability.
@@
Two-phase Continual Fine-tuning Model Phase 1 (P1) Phase 2 (P2) MLQA $(\uparrow)$ XLSUM $(\uparrow)$ XQuAD $(\uparrow)$ GMMLU $(\uparrow)$ Average Dataset Dataset P1 P2 P1 P2 P1 P2 P1 P2 P1 P2 Mistral-7B OpenOrca mOpenOrca 0.435 0.360 0.007 0.008 0.556 0.643 0.433 0.312 0.358 0.331 Instruct mOpenOrca 0.246 0.155 0.012 0.040 0.351 0.323 0.440 0.279 0.262 0.20 LLaMA-3-8B OpenOrca mOpenOrca 0.401 0.453 0.017 0.006 0.499 0.531 0.242 0.513 0.290 0.376 Instruct mOpenOrca 0.609 0.604 0.048 0.048 0.712 0.713 0.250 0.233 0.405 0.40 Dataset Mixture Model Dataset Mixture MLQA $(\uparrow)$ XLSUM $(\uparrow)$ XQuAD $(\uparrow)$ GMMLU $(\uparrow)$ Average Mistral-7B OpenOrca mOpenOrca 0.201 0.128 0.071 0.277 0.169 LLaMA-3-8B OpenOrca mOpenOrca 0.224 0.034 0.091 0.364 0.178
@@
Table B3: Multilingual Ability results for two-phase Continual Fine-tuning (CFT). With green, we highlight an increase in a model’s Multilingual ability post Phase 2 fine-tuning. Likewise, red highlights a decline in a model’s Multilingual ability.
English Ability.
@@
Table B2 present the english ability numbers of our ablations on the OpenOrca-mOpenOrcaand Instruct-mOpenOrcadatasets using Mistral-7B and LLaMA-3-8B models. When the datasets are pairwise not similar, i.e., Instruct-mOpenOrca, Mistral-7B shows a significant decline in the average English ability, from 0.450 in Phase 1 to 0.30 in Phase 2. Likewise, LLaMA-3-8B also experiences a decrease, dropping from 0.372 to 0.138 on average.
@@
In contrast, when the pairwise datasets are similar, i.e., OpenOrca and mOpenOrca, Mistral-7B sees a marginal increase between the phases $(0.407\rightarrow 0.413)$, on average. LLaMA-3-8B’s performance sees an improvement in the average English ability, from 0.337 to 0.356.
Multilingual Ability.
@@
Table B3 tabulates the results for multilingual ability. We see an improvement in the average multilingual ability for the OpenOrca-mOpenOrca dataset pair for LLaMA-3-8B. For Mistral-7B, there is a marginal drop $(0.358\to 0.331)$. For Instruct-mOpenOrca, with LLaMA-3-8B, the average multilingual ability is virtually the same across tasks (0.405 vs. 0.40). However, for Mistral-7B, we see a slight drop in the average language ability, driven primarily due to a decline in performance for MLQA.
Model Phase 1 (P1) Phase 2 (P2) IFEval $(\uparrow)$ Alpaca Eval $(\uparrow)$ MMLU $(\uparrow)$ HellaSwag $(\uparrow)$ Average Dataset Dataset P1 P2 P1 P2 P1 P2 P1 P2 P1 P2 Mistral-7B MultiAlpaca Alpaca 0.245 0.290 0.120 0.114 0.528 0.430 0.476 0.510 0.342 0.336 LLaMA-3-8B MultiAlpaca Alpaca 0.245 0.340 0.038 0.065 0.570 0.540 0.577 0.590 0.357 0.384 Mistral-7B mOpenOrca OpenOrca 0.190 0.310 0.091 0.055 0.410 0.490 0.520 0.510 0.303 0.341 LLaMA-3-8B mOpenOrca OpenOrca 0.314 0.340 0.0 0.0 0.530 0.540 0.522 0.590 0.342 0.368
Table B4: English Ability results for two-phase Continual Fine-tuning (CFT)
Model Phase 1 Phase 2 MLQA $(\uparrow)$ XLSUM $(\uparrow)$ XQuAD $(\uparrow)$ Average Dataset Dataset Phase 1 Phase 2 Phase 1 Phase 2 Phase 1 Phase 2 Phase 1 Phase 2 Mistral-7B MultiAlpaca Alpaca 0.122 0.230 0.021 0.030 0.122 0.090 0.088 0.116 LLaMA-3-8B MultiAlpaca Alpaca 0.363 0.340 0.048 0.040 0.058 0.030 0.157 0.134 Mistral-7B mOpenOrca OpenOrca 0.165 0.160 0.077 0.070 0.140 0.180 0.127 0.137 LLaMA-3-8B mOpenOrca OpenOrca 0.057 0.0 0.038 0.0 0.047 0.0 0.047 0.0
Table B5: Multilingual Ability results for two-phase Continual Fine-tuning (CFT)
@@
Appendix C Reverse Order CFT Result Analysis
In Tables B4 and B5, we reverse the order of Phase 1 and Phase 2 datasets, first fine-tuning on the multilingual dataset followed by its English counterpart. For Mistral-7B (MultiAlpaca-Alpaca), the average performance is 0.226, and for LLaMA-3-8B (MultiAlpaca-Alpaca), it is 0.259. Compared to the mixture setting and the Alpaca-MultiAlpaca configuration (§4), we observe that English ability benefits from multilingual fine-tuning in Phase 1, resulting in performance comparable to the data mixture setting. However, fine-tuning on English data in Phase 2 leads to a drastic drop in multilingual ability, yielding worse results than both the mixture setting and the two-phase setup considered in the main paper.
Appendix D Mitigating Strategies